# Towards Next-Generation Recommender Systems: A Benchmark for Personalized Recommendation Assistant with LLMs

### v1

@@ Appendix G Prompts for Model Testing
Appendix H Prompts for Query Construction
The prompts we use for constructing queries with LLMs are shown in Table LABEL:tab:query_generation_prompt.
Table 6. Summarization of prompts for query generation
Task Prompt
Explicit Condition Query You are given a set of attributes. Please simulate a real user and generate a natural language query covering these attributes to search or request recommendations for related movies. The attributes are {Explicit Conditions} …
Implicit Condition Query You are given a set of attributes and relevant information for [MASK] attributes. Your task is to generate a query that meets the following criteria: - Query should ask for items that share the input attributes. - Do not directly mention the [MASK] attribute that has additional relevant information (e.g., name of cinematography or starring role) in queries. Instead, describe the [MASK] attribute using the relevant information provided. The input attributes and relevant information is: …
Misinformed Condition Query You are given a set of attributes and relevant information for one of the attributes. Your task is to generate query that meets the following criteria: - Imagine you are a real user, queries should ask for movies that share the input attributes. - One of the attributes is provided with relevant information, such as the movie name, the relation between the movie and the attribute, and the person’s name. You should describe this attribute using the relevant information provided. - There may be factual errors between the relevant information and the corresponding attribute, but you still need to describe it based on the relevant information provided without making corrections or explanations. The input attributes and relevant information is: …
Interest-based Query ## Input Introduction You are given a “Popular Movie List” and “Previous Movie Statistics”. “Popular Movie List” represents movies collectively watched by multiple users, while ”Previous Movie Statistics” refers to the statistics of movies that some users watched before watching the “Popular Movie List”. Here is an example: - Popular Movie List: [A, D, M, …] - Previous Movie Statistics: - [Z, V, K], count: 6 - [M, O, Z,Y], count: 5 - [Y, C, E, O, Z], count: In this example, 6 users watched [A, D, M, …] after watching [Z, V, K] ## Task Introduction 1. **Step 1: Generate Reasons** Please use all your knowledge about these movies to analyze why users watch movies in the Popular Movie List after watching Previous Movies. What is the relationship between them? Do not give reasons that are too far-fetched or broad. If you think it cannot be explained, don’t force an explanation. 2. **Step 2: Generate User Queries and Answers** For each reason you generate: - Create **realistic user queries**. These queries should simulate how a real user might ask recommender systems for movie recommendations using natural language. The query should be as complex and rich as possible, close to the reason, and specific enough not to make the recommendation system recommend movies other than the subset of answers. - Provide a **subset of the Popular Movie List/Previous Movies** as the answer, ensuring the recommendations align with the reason and query. This subset will serve as the answer to the query, so do not mention the movie title of the subset in the query. ## Output Format The output must follow this JSON structured format: [ ‘reason ’: , ‘query ’: , ‘movie subset ’: , ‘reason ’: , ‘query ’: , ‘movie subset ’: , … ]
Demographics-based Query You are tasked with analyzing a specific user group and their movie preferences. Based on the provided data, you should generate structured reasons, user queries, and answers in a clear and organized format. Follow these steps to complete the task: ## **Input Data**: 1. **User Group**: A description of the user group. 2. **List of Movies**: A ranked list of movies relevant to the user group, including their titles, release years, TF-IDF scores, and the percentage of users in the group who viewed them. ## **Your Task**: 1. **Step 1: Generate Reasons** Analyze the user group description and the provided movie list. Generate a structured list of reasons explaining why this user group prefers these movies. 2. **Step 2: Generate User Queries and Answers** For each reason you generate: - Create **realistic user queries** that naturally combine the reason and user group characteristics. These queries should simulate how a real user from the group might ask recommender systems for movie recommendations using natural language. - Provide a **subset of the movie list** as the answer, ensuring the recommendations align with the reason and query. Use the movie title to represent a movie.

### v2

@@ Appendix A Prompts for Model Testing
Appendix B Case Study
We present a case study to illustrate how LLMs handle complex user queries. In this example, the user seeks movies with the same cinematographer as two given films. The model DeepSeek-V3, lacking strong reasoning abilities, mistakenly lists movies by Scott Ambrozy, leading to incorrect recommendations. In contrast, DeepSeek-R1 first verifies whether Scott Ambrozy is the cinematographer, then correctly identifies Owen Roizman as the actual cinematographer. It finally infers that the user intended to ask about Roizman’s films, providing accurate recommendations.
Case Study Query: I recently enjoyed the cinematography in Taps (1981) and Absence of Malice (1981), both of which showcased Scott Ambrozy’s incredible work. I’m eager to find more films that feature his cinematography style. Any recommendations? Ground Truth: 3 Days of the Condor (a.k.a. Three Days of the Condor) (1975), Network (1976), The Exorcist (1973), The Stepford Wives (1975) ; or ”None” Output (DeepSeek-V3): The Chosen (1981), Four Friends (1981), Split Image (1982), Heart Like a Wheel (1983) Output (DeepSeek-R1): The French Connection (1971), The Exorcist (1973), Network (1976), Three Days of the Condor (1975), The Electric Horseman (1979), Straight Time (1978), The Prince of the City (1981), Wyatt Earp (1994). Note: The films listed above feature the cinematography of Owen Roizman, who shot Taps and Absence of Malice. Scott Ambrozy appears to be a confusion or typo in the query. Reasoning (DeepSeek-R1): … Wait, first, who is Scott Ambrozy? I know some big cinematographers, but this name isn’t ringing a bell. Maybe I should verify if there’s a Scott Ambrozy in the industry. Let me think. Maybe the user misspelled the name? Looking up Taps on IMDb, yeah, Owen Roizman is the cinematographer for Taps. Absence of Malice also lists Owen Roizman. So Scott Ambrozy might be a confusion with another name. So the user probably meant Owen Roizman. So to answer correctly, I should recommend other films by Owen Roizman. Since the user wants movies with his cinematography style, I should list those. Roizman’s notable works include…
Ethical Considerations
Our work utilizes an LLM and public recommendation datasets for query generation. We have described the details of the generation process of our dataset in the main text of our paper. It is worth noting that all data in our dataset has been checked by human annotators to avoid harmful content. Therefore, our dataset does not pose any potential harm to human society.